# Unit-generated orders of real quadratic fields I. Class number bounds

### v2

@@ 4. Unit-generated orders with trivial class group
We study 1.4, the class number one problem for unit-generated quadratic orders. We give a complete solution for maximal orders. For the general case, we give a table of known solutions, conjectured to be complete.
@@ 4.1. Unit-generated maximal orders with class number one
We prove 1.5, by analysis of cases.
@@ 4.1.1. $\Delta^{-}_{n}=n^{2}+4$
    If $n\equiv 2\,\,\left(\operatorname{mod}\,4\right)$ (regardless of whether $\mathcal{O}_{\Delta}$ is maximal or nonmaximal), the class number $h_{\Delta}$ is even.
To prove the result, we use known criteria for oddness of class numbers of quadratic orders.
@@ 4.1.2. $\Delta^{+}_{n}=n^{2}-4$
In 1987 Mollin [38] conjectured that $h_{n^{2}-4}>1$ when $\Delta=n^{2}-4$ is squarefree and $n\geq 21$. This case is resolved by a result of 2007 Byeon, Kim, and Lee [10, Theorem 1.2], who solved the class number $1$ problem for squarefree discriminants $\Delta=n^{2}-4$, showing all solutions have $n<21$. Their proof uses the general method of Biró. For $\Delta=n^{2}-4$ the squarefree assumption is equivalent to $\Delta$ being the discriminant of a maximal order with $n$ odd.
@@
The solution list in 4.4 was determined numerically using the bound $n\leq 21$.
We resolve the fourth subcase, of orders having fundamental discriminants $\Delta=n^{2}-4$ having $n$ even, using genus theory.
@@
    If $\mathcal{O}_{\Delta}$ is a maximal order with $n=2m$ even, and with class number $h_{\Delta}=1$, then $m\in\{0,2\}$, giving $n\in\{0,4\}$ with discriminant $\Delta\in\{-4,12\}$.
@@
    If $n\equiv 2\,\,\left(\operatorname{mod}\,4\right)$ (regardless of whether $\mathcal{O}_{\Delta}$ is maximal or nonmaximal), and $m>3$ then the class number $h_{\Delta}$ is even. For $m=3$, giving $\Delta=32$, we have $h_{\Delta}=1$.
@@
(1) We first treat small $m$. First, $\Delta=-4$ coming from $m=0$,$n=0$, is a fundamental discriminant, and has $h_{\Delta}=1$. Next $\Delta=12$ coming from $m=2$, $n=4$, is a fundamental discriminant, and has $h_{\Delta}=1$. These give the two solutions listed in (1). Next $\Delta=0$, coming from $m=1$, $n=2$, is ruled out as not being a quadratic discriminant Finally $\Delta=32$, coming from $m=3$, $n=6$, is a non-fundamental discriminant, and has $h_{\Delta}=1$, giving the discriminant listed in (2).
@@
Then $\Delta=4(4k^{2}-1)=4(2k+1)(2k-1)$. Let $g=\gcd(2k+1,2k-1)$. Then $g$ is odd and g| 2g\,|\,2 so $g=1$. Now $2k+1\geq 5$ has an odd prime factor, call it $p$, and $2k-1\geq 3$ has an odd prime factor $q$ distinct from $p$, by relative primality. Thus $4pq\,|\,\Delta$.
If $\Delta$ is a fundamental discriminant, the condition $4pq\,|\,\Delta$ is enough to force $h_{\Delta}$ to be even, by checking the criterion (3) of 4.3. Here conditions $(a),(b)$ are ruled out since $pq\,|\,\Delta$, condition $(c)$ is ruled out since $4\,|\,\Delta$, and $(d)$ is ruled out since $\Delta\geq 60$. Thus there are no such fundamental $\Delta$ with $h_{\Delta}=1$ in Case 1.
@@
Now $\Delta=4(m^{2}-1)$. Since $m$ is odd $x^{2}\equiv 1\,\,\left(\operatorname{mod}\,8\right)$, so $32\,|\,\Delta$, and $\Delta$ is a quadratic discriminant. Now the class number $h_{\Delta}$ is even (whether $\Delta$ is fundamental or non-fundamental) by checking the criterion (2) of 4.3 for wide class number being odd. Indeed $32\,|\,\Delta$ rules out conditions (a)–(d) and $\Delta\geq 60$ rules out (e). So there are no $\Delta$, with $h_{\Delta}=1$, fundamental or non-fundamental, in Case 2.

### v3

@@ 4. Unit-generated orders with trivial class group
We classify all unit-generated quadratic orders with class number one.
@@ 4.1. Unit-generated maximal orders with class number one
We prove 1.3 by analysis of cases.
@@ 4.1.1. $\Delta^{-}_{n}=n^{2}+4$
    If $n\equiv 2\,\,\left(\operatorname{mod}\,4\right)$, then the class number $h_{\Delta}$ is even, regardless of whether $\mathcal{O}_{\Delta}$ is maximal or non-maximal.
To prove the result, we use known criteria from genus theory for oddness of class numbers of quadratic orders.
@@ 4.1.2. $\Delta^{+}_{n}=n^{2}-4$
We divide the case $\Delta=\Delta^{+}=n^{2}-4$ ($n\geq 0$) into two subcases according to the parity of $n$. For $n$ odd, in 1987 Mollin [38] conjectured that $h_{n^{2}-4}>1$ when $\Delta=n^{2}-4$ is squarefree and $n\geq 21$. For $\Delta=n^{2}-4$ the squarefree assumption is equivalent to $\Delta$ being the discriminant of a maximal order with $n$ odd. Mollin’s conjecture was proved in 2007 by Byeon, Kim, and Lee [10, Theorem 1.2]. Their proof uses an adaptation of the general method of Biró.
@@
We resolve the remaining subcase, of orders having fundamental discriminants $\Delta=n^{2}-4$ having $n$ even, using genus theory.
@@
    If $n=2m$ is even, and $\mathcal{O}_{\Delta}$ is a maximal order with class number $h_{\Delta}=1$, then $n\in\{0,4\}$ with discriminant $\Delta\in\{-4,12\}$.
@@
    If $n=2m\equiv 2\,\,\left(\operatorname{mod}\,4\right)$, and $m\geq 3$, then the class number $h_{\Delta}$ is even, regardless of whether $\mathcal{O}_{\Delta}$ is maximal or non-maximal. For $m=3$ with $n=6$, we have $\Delta=32$ and $h_{\Delta}=1$.
@@
(1) We first treat small $m$. First, $\Delta=-4$, coming from $(m,n)=(0,0)$, is a fundamental discriminant and has $h_{\Delta}=1$. Next, $\Delta=12$, coming from $(m,n)=(2,4)$, is a fundamental discriminant and has $h_{\Delta}=1$. These give the two solutions listed in (1). Next, $\Delta=0$, coming from $(m,n)=(1,2)$, is ruled out as not being a quadratic discriminant. Finally, $\Delta=32$, coming from $(m,n)=(3,6)$, is a non-fundamental discriminant and has $h_{\Delta}=1$, giving the discriminant listed in (2).
@@
In this case we consider only fundamental discriminants. Then $\Delta=4(4k^{2}-1)=\mbox{$4(2k+1)(2k-1)$}$. Let $g=\gcd(2k+1,2k-1)$. Then $g$ is odd and g| 2g\,|\,2, so $g=1$. Now $2k+1\geq 5$ has an odd prime factor, call it $p$, and $2k-1\geq 3$ has an odd prime factor $q$ distinct from $p$, by relative primality. Thus $4pq\,|\,\Delta$.
We show that $2\,|\,h_{\Delta}$ when $\Delta$ is a fundamental discriminant, by excluding all cases of odd class number in criterion (3) of 4.3. Conditions $(a),(b)$ are ruled out since $pq\,|\,\Delta$, condition $(c)$ is ruled out since $4\,|\,\Delta$, and condition $(d)$ is ruled out since $\Delta\geq 60$. Case 1 follows.
@@
In this case we consider all real quadratic discriminants $\Delta=4(m^{2}-1)$. Since $m$ is odd, $x^{2}\equiv 1\,\,\left(\operatorname{mod}\,8\right)$, so $32\,|\,\Delta$, and $\Delta$ is a quadratic discriminant. Now the class number $h_{\Delta}$ is even (whether $\Delta$ is fundamental or non-fundamental) by checking the criterion (2) of 4.3 for wide class number being odd. Indeed $32\,|\,\Delta$ rules out conditions (a)–(d), and case (e) is ruled out since $\Delta\geq 60$. Case 2 follows.
@@
For narrow Richaud–Degert maximal orders of class number one, the remaining case not covered by the results above consists of all squarefree discriminants $\Delta=4n^{2}+1$, coming from case (2) of 2.5). This case is Chowla’s conjecture, which was formulated in 1976 (originally only for prime $\Delta$) in [13, p. 48]. It was settled by Biró [8, p. 179] in 2003 using his method.
Theorem 4.6 (Biró).
If $\Delta=4n^{2}+1$ is the discriminant of a maximal order having class number $h_{\Delta}=1$, then necessarily $n\leq 1861$. The solutions are $n\in\{1,2,3,5,7,13\},$ with discriminants $\Delta\in\{5,17,37,101,197,677\}.$
The condition that $\Delta=4n^{2}+1$ is the discriminant of a maximal order is equivalent to the requirement that $4n^{2}+1$ be squarefree, used in Biró’s formulation of the result. The solution list in 4.6 was determined numerically using the bound $n\leq 1861$.
4.6 will be used in the non-maximal order case treated in Section 4.2.
4.2. Unit-generated non-maximal orders with class number one
We classify all non-maximal unit-generated quadratic orders $\mathcal{O}$ of class number one. For such orders $\mathcal{O}$, all their over-orders, including the maximal order, necessarily have class number one. We must show an additional key property: If a non-maximal quadratic order $\mathcal{O}$ has class number one and is unit-generated, then all its over-orders $\mathcal{O}^{\prime}$ necessarily are also unit-generated, with one exceptional case that is “almost” unit-generated.
We handle the case of prime relative conductor $\left[\mathcal{O}^{\prime}:\mathcal{O}\right]=p$ first.
Proposition 4.7.
Suppose that $\Delta=p^{2}\Delta^{\prime}$, where $\Delta$ and $\Delta^{\prime}$ are non-square discriminants and $p$ is a prime number. If $\mathcal{O}_{\Delta}$ is a unit-generated order and $h_{\Delta}=h_{\Delta^{\prime}}$, then one of the following is true:
  • (1)
    $\mathcal{O}_{\Delta^{\prime}}$ is a unit-generated order and $p<\frac{\log(p\sqrt{\Delta^{\prime}}+1)}{\log(\sqrt{\Delta^{\prime}}-1)}+1$, from which it follows that $p\leq 17$.
  • (2)
    $p=2$ and $\Delta^{\prime}\equiv 1\,\,\left(\operatorname{mod}\,8\right)$.
Proof.
Note that the quotient ideal $\left(\mathcal{O}_{\Delta}:\mathcal{O}_{\Delta^{\prime}}\right)=p\mathcal{O}_{\Delta^{\prime}}$. We have by [35, Thm. 6.5] the exact sequence
$1\to\frac{\mathcal{O}_{\Delta^{\prime}}^{\times}}{\mathcal{O}_{\Delta}^{\times}}\xrightarrow{\iota}\frac{\left(\mathcal{O}_{\Delta^{\prime}}/p\mathcal{O}_{\Delta^{\prime}}\right)^{\times}}{\left(\mathcal{O}_{\Delta}/p\mathcal{O}_{\Delta^{\prime}}\right)^{\times}}\xrightarrow{\phi}\operatorname{Cl}(\mathcal{O}_{\Delta})\xrightarrow{\psi}\operatorname{Cl}(\mathcal{O}_{\Delta^{\prime}})\to 1.$ (4.1)
The equality $h_{\Delta}=h_{\Delta^{\prime}}$ implies that $\psi$ is an isomorphism, so $\phi=0$ and $\iota$ is an isomorphism. There is a ring isomorphism $\mathcal{O}_{\Delta^{\prime}}/p\mathcal{O}_{\Delta^{\prime}}\cong\mathbb{Z}[x]/(p,g(x))\cong\mathbb{F}_{p}[x]/(g(x))$ where $g(x)=x^{2}-\Delta^{\prime}x+\frac{(\Delta^{\prime})^{2}-\Delta^{\prime}}{4}$ having $\operatorname{disc}(g)=\Delta^{\prime}$. Set $\omega=\frac{\Delta^{\prime}+\sqrt{\Delta^{\prime}}}{2}$. The subring $\mathcal{O}_{\Delta}/p\mathcal{O}_{\Delta^{\prime}}$ then has a compatible isomorphism $\mathcal{O}_{\Delta}/p\mathcal{O}_{\Delta^{\prime}}=\frac{p\omega\mathbb{Z}+\mathbb{Z}}{p\omega\mathbb{Z}+p\mathbb{Z}}\cong\mathbb{F}_{p}$. The quotient of unit groups is determined by the Kronecker symbol $\left(\frac{\Delta^{\prime}}{p}\right)$:
$\frac{\left(\mathcal{O}_{\Delta^{\prime}}/p\mathcal{O}_{\Delta^{\prime}}\right)^{\times}}{\left(\mathcal{O}_{\Delta}/p\mathcal{O}_{\Delta^{\prime}}\right)^{\times}}\cong\frac{\left(\mathbb{F}_{p}[x]/(g(x))\right)^{\times}}{\mathbb{F}_{p}^{\times}}\cong\begin{cases}\mathbb{Z}/(p-1)\mathbb{Z}&\text{if }\left(\frac{\Delta^{\prime}}{p}\right)=1,\\ \mathbb{Z}/p\mathbb{Z}&\text{if }\left(\frac{\Delta^{\prime}}{p}\right)=0,\\ \mathbb{Z}/(p+1)\mathbb{Z}&\text{if }\left(\frac{\Delta^{\prime}}{p}\right)=-1.\end{cases}$ (4.2)
It follows from the fact that $\iota$ in (4.1) is an isomorphism that the index
$\mathcal{O}_{\Delta^{\prime}}^{\times}:\mathcal{O}_{\Delta}^{\times}]=p-\left(\frac{\Delta^{\prime}}{p}\right).$ (4.3)
If this index is not $1$, then $\mathcal{O}_{\Delta^{\prime}}$ contains a unit that is not in $\mathcal{O}_{\Delta}$, so since $\mathcal{O}_{\Delta}$ is unit-generated and there are no other orders between $\mathcal{O}_{\Delta^{\prime}}$ and $\mathcal{O}_{\Delta}$, we conclude that $\mathcal{O}_{\Delta^{\prime}}$ is unit-generated. If the index is $1$, then $p=2$ and $\left(\frac{\Delta^{\prime}}{p}\right)=1$, that is, $\Delta^{\prime}\equiv\pm 1\,\,\left(\operatorname{mod}\,8\right)$, so $\Delta^{\prime}\equiv 1\,\,\left(\operatorname{mod}\,8\right)$ because it is a discriminant.
It remains to show the upper bound on $p$ in case (1). Let $u>1$ be a generator for $\mathcal{O}_{\Delta^{\prime}}^{\times}/\{\pm 1\}$ and let $v>1$ be a generator for $\mathcal{O}_{\Delta}^{\times}/\{\pm 1\}$. Let $\epsilon=\left(\frac{\Delta^{\prime}}{p}\right)$. By (4.3), $v=u^{p-\epsilon}$. Moreover,
$\sqrt{\Delta^{\prime}}-1=u-\sigma(u)-1$ $<u<u-\sigma(u)+1=\sqrt{\Delta^{\prime}}+1,$ (4.4)
$p\sqrt{\Delta^{\prime}}-1=v-\sigma(v)-1$ $<v<v-\sigma(v)+1=p\sqrt{\Delta^{\prime}}+1,$ (4.5)
for the nontrivial Galois automorphism $\sigma$. Thus,
$p$ $=\frac{\log v}{\log u}+\epsilon<\frac{\log(p\sqrt{\Delta^{\prime}}+1)}{\log(\sqrt{\Delta^{\prime}}-1)}+1.$ (4.6)
Finally, note that the function $f(x,y)=\frac{\log(xy+1)}{\log(y-1)}$ is a decreasing function of $y$ for any $x,y>1$, because
$\frac{\partial}{\partial y}f(x,y)$ $=\frac{xy\log(y-1)-x\log(y-1)-xy\log(xy+1)-\log(xy+1)}{(y-1)(xy+1)\left(\log(y-1)\right)^{2}}$ (4.7)
$<\frac{xy\log(xy+1)-x\log(y-1)-xy\log(xy+1)-\log(xy+1)}{(y-1)(xy+1)\left(\log(y-1)\right)^{2}}$ (4.8)
$=\frac{-x\log(y-1)-\log(xy+1)}{(y-1)(xy+1)\left(\log(y-1)\right)^{2}}<0.$ (4.9)
From (4.6), we can conclude that
$p<f(p,\sqrt{\Delta^{\prime}})+1\leq f(p,\sqrt{5})+1.$ (4.10)
By again taking a derivative, we can see that the function $g(x)=f(x,\sqrt{5})+1-x$ is decreasing for $x>4.3$; also, $g(19)\approx-0.2<0$. Thus, $g(p)>0$ (that is, (4.10)) implies $p\leq 17$. ∎
We now prove the key property for general relative conductor $\left[\mathcal{O}^{\prime}:\mathcal{O}\right]=f$.
Proposition 4.8.
Suppose that $\Delta=f^{2}\Delta^{\prime}$, where $\Delta$ and $\Delta^{\prime}$ are non-square discriminants. If $\mathcal{O}_{\Delta}$ is a unit-generated order and $h_{\Delta}=h_{\Delta^{\prime}}$, then one of the following is true:
  • (1)
    $\mathcal{O}_{\Delta^{\prime}}$ is a unit-generated order.
  • (2)
    $\mathcal{O}_{\Delta^{\prime}}$ is not a unit-generated order, but $\mathcal{O}_{4\Delta^{\prime}}$ is a unit-generated order, 2|f2\,|\,f, and $\Delta^{\prime}\equiv 1\,\,\left(\operatorname{mod}\,8\right)$.
Proof.
Let $e$ be the smallest divisor of $f$ such that $\mathcal{O}_{e^{2}\Delta^{\prime}}$ is unit-generated. If $e=1$, we are done; this is case (1).
Otherwise, suppose $e\geq 2$ so $\mathcal{O}_{\Delta^{\prime}}$ is not a unit-generated order. Let $p$ be a prime divisor of $e$, and choose $p$ to be odd unless $e$ is a power of two. We must have $h_{\Delta^{\prime}}\leq h_{(e/p)^{2}\Delta^{\prime}}\leq h_{e^{2}\Delta^{\prime}}\leq h_{\Delta}$, but also $h_{\Delta}=h_{\Delta^{\prime}}$, so in fact $h_{\Delta^{\prime}}=h_{(e/p)^{2}\Delta^{\prime}}=h_{e^{2}\Delta^{\prime}}=h_{\Delta}$. By 4.7 applied to the pair of discriminants $(e^{2}\Delta^{\prime},(e/p)^{2}\Delta^{\prime})$, we conclude that either:
  • (1’)
    $\mathcal{O}_{(e/p)^{2}\Delta^{\prime}}$ is a unit generated order, or
  • (2’)
    $p=2$ and $(e/p)^{2}\Delta^{\prime}\equiv 1\,\,\left(\operatorname{mod}\,8\right)$.
The minimality of $e$ rules out case (1’). Thus, $p=2$, which forces $e$ to be a power of two, and $(e/p)^{2}\Delta^{\prime}\equiv 1\,\,\left(\operatorname{mod}\,8\right)$, which subsequently forces $e=2$ and $\Delta^{\prime}\equiv 1\,\,\left(\operatorname{mod}\,8\right)$. We conclude that $\mathcal{O}_{4\Delta^{\prime}}$ is a unit-generated order, 2|f2\,|\,f, and $\Delta^{\prime}\equiv 1\,\,\left(\operatorname{mod}\,8\right)$. This is case (2). ∎
Proof of 1.4.
It may be checked directly that each of these (non-maximal) orders are unit-generated and have class number $1$. We focus on the proof of the converse, that every unit-generated real quadratic non-maximal order of class number $1$ is contained in the set $\{32,45,117\}\cup\{20,68,125\}$.
Suppose that $\mathcal{O}_{\Delta}$ is a unit-generated order and $h_{\Delta}=1$. Let $\Delta_{0}$ be the associated fundamental discriminant, and write $\Delta=f^{2}\Delta_{0}$, with $f>1$. We must have $h_{\Delta_{0}}=1$. By 4.8, one of the following holds.
  • (1)
    $\mathcal{O}_{\Delta_{0}}$ is a unit-generated order.
  • (2)
    $\mathcal{O}_{\Delta_{0}}$ is not a unit-generated order, but $\mathcal{O}_{4\Delta_{0}}$ is a unit-generated order, 2|f2\,|\,f, and $\Delta_{0}\equiv 1\,\,\left(\operatorname{mod}\,8\right)$.
In case (1), 1.3 implies that the fundamental discriminant $\Delta_{0}$ is contained in the list
$\Delta_{0}\in\{5,8,12,13,21,29,53,77,173,293,437\}.$ (4.11)
For each fundamental $\Delta_{0}$ in this list, our results allow us to enumerate all possible non-fundamental values for $\Delta$ by a finite computation described as follows.
  1. (S1)
    Check whether $\mathcal{O}_{4\Delta_{0}}$ is a unit-generated order of class number $1$. (This happens for $\Delta_{0}\in\{5,8\}$, giving $\Delta\in\{20,32\}$.)
  2. (S2)
    For those cases where (S1) holds, also check whether $\mathcal{O}_{4p^{2}\Delta_{0}}$ is a unit-generated order of class number $1$ for each of the finite list of primes satisfying $p\leq 17$. (This never happens.)
  3. (S3)
    Thus, by 4.7, there do not exist any primes $p$ such that $\mathcal{O}_{4p^{2}\Delta_{0}}$ is a unit-generated order of class number $1$.
  4. (S4)
    If $\Delta$ is even and $\Delta\neq 4\Delta_{0}$, applying the criterion of 4.8 to a pair of the form $(\Delta,4p^{2}\Delta_{0})$ rules out $\Delta$.
  5. (S5)
    Check whether $\mathcal{O}_{p^{2}\Delta_{0}}$ is a unit-generated order of class number $1$ for each of the finite list of odd primes satisfying $p\leq 17$. (This happens for $(\Delta_{0},p)\in\{(5,3),(13,3),(5,5)\}$, giving $\Delta\in\{45,117,125\}$.)
  6. (S6)
    For those cases where (S5) holds, also check whether $\mathcal{O}_{p^{2}q^{2}\Delta_{0}}$ is a unit-generated order of class number $1$ for each of the finite list of odd primes satisfying $q\leq 17$, allowing $q=p$. (This never happens.)
  7. (S7)
    Thus, by 4.7, there do not exist any primes $p,q$ such that $\mathcal{O}_{p^{2}\Delta_{0}}$ or $\mathcal{O}_{p^{2}q^{2}\Delta_{0}}$ is a unit-generated order of class number $1$, except for those found in (S5).
  8. (S8)
    If $\Delta$ is odd, $\Delta\neq\Delta_{0}$, and $\Delta$ was not found in (S5), applying 4.8 to a pair of the form $(\Delta,p^{2}\Delta_{0})$ or $(\Delta,p^{2}q^{2}\Delta_{0})$ to rule out $\Delta$. In the case 4.8(2), we are using the conclusion of (S4) to say that $\mathcal{O}_{4p^{2}\Delta_{0}}$ and $\mathcal{O}_{4p^{2}q^{2}\Delta_{0}}$ cannot be unit-generated orders of class number $1$.
This computation yields for $\Delta=n^{2}-4$ with $n\in\{6,7,11\}$ the non-maximal discriminants $\Delta\in\{32,45,117\}$. It yields for $\Delta=n^{2}+4$ with $n\in\{4,11\}$ the non-maximal discriminants $\Delta\in\{20,125\}$.
In case (2), $\Delta_{0}\equiv 1\,\,\left(\operatorname{mod}\,8\right)$, $\mathcal{O}_{\Delta_{0}}$ is not a unit-generated order, and $4\Delta_{0}=n^{2}\pm 4$ for some integer $n$. Write $n=2m$. If $4\Delta_{0}=n^{2}-4$, then $\Delta_{0}=m^{2}-1\equiv 0,3,7\,\,\left(\operatorname{mod}\,8\right)$, a contradiction. Thus $4\Delta_{0}=n^{2}+4$, so $\Delta_{0}=m^{2}+1$. Necessarily $m$ must be even, say $m=2k$, so $\Delta_{0}=4k^{2}+1$. Since $\Delta_{0}$ is fundamental and $h(\Delta_{0})=1$, 4.6 applies to give $k\in\{1,2,3,5,7,13\}$, so
$\Delta_{0}\in\{5,17,37,101,197,677\}.$ (4.12)
Only $\Delta_{0}=17$ satisfies $\Delta_{0}\equiv 1\,\,\left(\operatorname{mod}\,8\right)$. It yields for $n=8$ the new non-maximal discriminant $\Delta=68$. There are no more, as one rules out all $\Delta=4p^{2}\Delta_{0}$ for $p\leq 17$ being unit-generated orders of class number one. ∎